# PCoreSet: Effective Active Learning through Knowledge Distillation from Vision-Language Models

### v1

@@ 4.2 Experimental Results
To assess whether PCoreSet effectively promotes structured bias propagation, we visualize the same metric used in Fig. 4(c) across AL rounds. We include an additional baseline, PCoreSet+R(Reverse), which inhibits coverage over the probability space by replacing $\arg\max$ with $\arg\min$ in line 4 of Alg. 2. As shown in Fig. 9, PCoreSet consistently yields the lowest KNN loss and highest cluster purity among all selection strategies. This confirms that PCoreSet promotes structured bias propagation as intended, resulting in improved AL performance. In contrast, PCoreSet+R exhibits the opposite behavior, supporting our claim that PCoreSet encourages bias propagation by maximizing sample diversity in the probability space.
@@
Figure 11: The heatmap of output probability vectors from different selection strategies in the first active learning round using the DTD dataset. See Sec. E.1 for results on other datasets.
@@
To evaluate the effectiveness of each active selection criterion, we analyze the progression of four strategies designed to be maximized throughout the active learning (AL) rounds: 1) Uncertainty [35], 2) Class balance [6], 3) Feature space diversity [76], and 4) Probability space diversity (PCoreSet). Criteria 1) and 2) are measured using normalized Shannon entropy, while 3) and 4) are evaluated using normalized average minimum distances. We report the average of each metric across 10 datasets, excluding ImageNet. As shown in Fig. 10, each method performs best on its respective objective: Entropy maximizes uncertainty, ClassBalanced optimizes the entropy of the class distribution, and Coreset maximizes feature space diversity. As expected, our PCoreSet method achieves the highest diversity in probability space, while also ranking third in uncertainty and class balance, and second in feature diversity. In contrast, traditional Coreset approaches [76], though effective in covering the feature space and offering moderate probabilistic diversity, perform poorly in class balance—showing their limitations under ActiveKD settings.
Figure 12: Images selected by PCoreSet on the Caltech101 dataset.
@@
To visualize the sample diversity achieved by PCoreSet, we present probability heatmaps of both the initially labeled samples and those selected by various active learning strategies after the first round, as shown in Fig. 11. The initial labeled set (1-shot per class) generally leads to correct classifications based on ground truth, though the probability distributions deviate from ideal values due to the influence of the teacher distillation signal. The heatmaps clearly show that PCoreSet selects samples with highly diverse probabilistic characteristics, effectively maximizing distributional coverage compared to the initial labeled set. We further visualize the selected images from the Caltech101 dataset in Fig. 12. PCoreSet selects not only visually distinct examples but also samples from different classes. Visualizations for other datasets are provided in Sec. E.2.
@@ 5 Conclusion, Limitation, and Future Work
We introduced ActiveKD, integrating active learning with knowledge distillation by leveraging VLMs’ zero- and few-shot capabilities. We discovered VLMs exhibit structured prediction bias that benefits student learning, and proposed Probabilistic Coreset (PCoreSet) to maximize diversity in probability space when selecting samples. Experiments across 11 datasets show PCoreSet outperforms existing methods, especially when labeled data is scarce, enabling efficient knowledge transfer under limited annotation budgets. Our work is currently limited to visual recognition with CLIP. Future research could extend ActiveKD to more complex tasks like object detection and segmentation, and incorporate instruction-tuned generalist models like FLAN [87] or LLaVA [55] to enhance effectiveness across diverse applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
  • Theoretical Analysis (Appendix A): provides formal analysis of bias propagation in knowledge distillation, including optimal student prediction analysis (Sec. A.1), formal characterization of teacher prediction bias (Sec. A.2), and mathematical proof of how biases transfer from teacher to student models (Prop. 2).
  • Implementation (Appendix B): detailed experimental setup including model configurations, training parameters.
  • Datasets (Appendix C): comprehensive description of the datasets used in our experiments, including ImageNet and other benchmarks used for evaluating bias propagation, as well as detailed analysis of class distributions across all datasets.
  • Additional Experiments (Appendix D): supplementary experimental results including evaluations on alternative architectures (e.g., MobileNetV2 in Sec. D.1), active learning without knowledge distillation, and few-shot teacher distillation scenarios.
  • Additional Qualitative Results (Appendix E): additional results of the proposed method, including heatmap visualizations of selected samples across datasets (Sec. E.1) and qualitative analysis of samples selected by our PCoreSet method (Sec. E.2).
@@ Appendix A Theoretical Analysis
This section presents a formal analysis of bias propagation in knowledge distillation, examining how teacher model biases influence student model predictions. We first analyze the optimal student prediction after distillation from a teacher model in Section A.1, building upon the analytical framework established in DHO [41]. Subsequently, in Section A.2, we formally characterize the mechanisms through which teacher prediction biases propagate to and constrain the distilled student model, providing the theoretical foundation for our debiasing strategies.
@@ A.1 Optimal Student Prediction after Distillation
Let us first establish our notation. We consider a student model with a feature extractor $h:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that maps inputs to feature representations, and a classification head $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ that maps features to logits, with the final prediction obtained by applying the softmax function $\sigma:\mathbb{R}^{C}\rightarrow\Delta^{C-1}$ to these logits, resulting in $f_{r}(x)=\sigma(g(h(x)))$, where $r$ denotes the active learning round. Following DHO [41], we consider two target probability distributions: the ground truth label distribution $y$, typically one-hot encoded vectors where $y_{c}=1$ for the true class $c$ and 0 elsewhere, and the teacher’s softened distribution $f(x)$ for input $x\in\mathcal{X}$.
@@
For the Dual-Head Approach, in the Dual-Head Optimization (DHO) framework [41], shared features $h(x)$ are extracted from input $x$ and two specialized classification heads are employed: $g_{\text{CE}}(h(x))$, optimized for ground truth labels using cross-entropy loss, and $g_{\text{KD}}(h(x))$, optimized for teacher predictions using KL divergence. This approach addresses the constrained optimization problem:
@@
The DHO framework [41] provides a formal proof that under appropriate optimization conditions, the dual-head inference is equivalent to single-head inference when setting $\alpha=\lambda$ and $\beta=\tau$, meaning that for any choice of distillation weight $\lambda$ and temperature $\tau$, the two approaches converge to the same prediction.
@@ A.2 Formal Definition of Teacher Prediction Bias
Let the teacher model $f$ exhibit structured prediction bias as defined in Def. 2, with $\{\mu_{k}\}_{k=1}^{K}$ and $\{r_{k}\}_{k=1}^{K}$. Assume the student model $f_{r}$ is trained via KD from $f$ using the loss $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, and satisfies $\sup_{x}\|f_{r}(x)-f^{*}(x)\|_{1}\leq\varepsilon$ for all $x\in\mathcal{X}$, where $f^{*}(x)=\lambda y+(1-\lambda)f(x)$ and $y\in\{0,1\}^{C}$ denotes the one-hot label of $x$. Then student predictions $f_{r}(x)$ also exhibit structured prediction bias as in Def. 2. Specifically, for every $x\in\mathcal{X}$, there exists $k\in[K]$ such that:
@@
where the propagated centroid is defined as $\hat{\mu}_{k}(x)=\lambda y+(1-\lambda)\mu_{k}$, and the adjusted radius is $\hat{r}_{k}=(1-\lambda)r_{k}+\varepsilon$. Since $y$ is one-hot and $\mu_{k}$ is fixed, the set of possible centers $\{\hat{\mu}_{k}(x)\}$ is finite and contained in $\Delta^{C-1}$, thus satisfying the condition of Def. 2 with at most $C\cdot K$ clusters.
@@ Appendix B Implementation Details
Algorithm 3 DHO Training with zero-shot CLIP [68] teacher
@@
2:$z\leftarrow g(x)$
3:$\hat{p}_{\text{CE}}\leftarrow\sigma(h_{\text{CE}}(z))$
4:$\hat{p}_{\text{KD}}\leftarrow\sigma(h_{\text{KD}}(z)/\beta)$
5:$\hat{p}\leftarrow\alpha\cdot\hat{p}_{\text{CE}}+(1-\alpha)\cdot\hat{p}_{\text{KD}}$
6:$\hat{y}\leftarrow\arg\max_{c}(\hat{p}_{c})$
@@
We employ the DHO [41] framework for knowledge distillation in our approach, as it provides a principled and effective mechanism for transferring knowledge from foundation teacher models to student models. DHO’s key advantage lies in its dual-head architecture that maintains separated optimization paths for labeled data and teacher knowledge, thereby preserving the distinct learning signals throughout training. Alg. 3 details the DHO training procedure with our zero-shot CLIP teacher, while Alg. 4 outlines the inference process that effectively combines both optimization pathways.
@@
Student: DINO self-supervised ResNet-50 [11] Teacher: CLIP ResNet-50 [68] Active learning rounds: 8 Initial setting: 1-shot (single image per class) Query size: 1,000 Unlabeled pool: 100,000 samples randomly selected with seed Random seeds: 5 different seeds KD parameters: $\zeta=0.01$, $\eta=2$, $\lambda=0.5$ DHO parameters: $\alpha=0.4$, $\beta=0.5$ Validation: No validation split used Epochs: 20 for first 7 rounds, 50 for final round Optimizer: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) Learning rate: $1\times 10^{-3}$ Weight decay: $1\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256)
@@
Student: ImageNet pre-trained ResNet-18 [29], MobileNetV2 [75], and ImageNet-21k [16] pre-trained ViT-Tiny [89] Teacher: CLIP ResNet-50 [68] Active learning rounds: 16 Initial setting: 1-shot (single image per class) Query size: Equal to number of classes per round Unlabeled pool: All training samples except labeled set Random seeds: 5 different seeds KD parameters: $\zeta=0.01$, $\eta=2$, $\lambda=0.5$ DHO parameters: $\alpha=0.5$, $\beta=1$ Validation: No validation split used Epochs: 200 for all rounds Optimizer: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) Learning rate: $1\times 10^{-3}$ Weight decay: $1\times 10^{-2}$ Batch size: 128 (labeled: 64, unlabeled: 64)
@@
Teacher: CLAP [77] on CLIP ResNet-50 [68] Setup: Following original CLAP paper [77] Modification: Learning rate reduced from 0.1 to 0.01 for better convergence
@@
Recognizing the potential for improving teacher network performance, particularly in low-data regimes, we explored few-shot learning techniques to enhance the distillation process. After considering various approaches [39, 103, 102, 42, 104, 43, 58, 100, 72, 98, 48], we incorporated CLAP [77], a method that notably does not require a validation set [77, 61, 60], as our teacher model. For training the teacher, we followed the setting from the original CLAP paper [77], with only one modification: reducing the learning rate from 0.1 to 0.01 for better convergence in our setting. This configuration represents a more realistic scenario wherein both teacher and student networks evolve simultaneously within the active distillation framework, better reflecting practical applications where pre-trained teacher models may not be available.
@@
BADGE [4] adopts a hybrid approach to maximize uncertainty and diversity simultaneously, utilizing the gradient of the final linear classifier layer which has $H\times C$ elements, where $H$ is the hidden dimension of the backbone feature and $C$ is the number of classes in the teacher model. However, DHO [41] extends beyond conventional single classifier architectures to dual classifiers that follow different optimization paths for labeled ground truth and teacher distillation signals. The dual classifier approach provides complementary information from both learning objectives, enhancing model performance through this specialized learning framework. In order to adopt DHO into the BADGE selection method, we manually calculate the gradients of both classifiers and concatenate them to form a gradient representation with $2\times H\times C$ elements. This concatenated gradient naturally extends the original BADGE selection algorithm while preserving its core k-means++ clustering mechanism.
@@
We implement the Class-Balanced Selection algorithm (Alg. 5) as our baseline method to address class distribution bias. The algorithm assigns pseudo-labels to unlabeled samples based on model predictions, analyzes class distribution in the labeled set, and computes inverse weights proportional to each class’s representation—giving higher weights to underrepresented classes. It then allocates the query budget across classes according to these weights and randomly selects samples from each class partition. This approach effectively counteracts class imbalance by prioritizing underrepresented classes.

### v2

@@ 4.3 Analysis
To assess whether PCoreSet promotes structured bias propagation, we track the same metric, i.e., KNN loss ($\ell_{2}$) and clster purity, used in Fig.˜4(c) across AL rounds. We also include a counterfactual baseline, PCoreSet+R (Reverse), which suppresses probability-space coverage by replacing $\arg\max$ with $\arg\min$ in line 4 of Algorithm˜2. As shown PCoreSet consistently achieves the lowest KNN loss and the highest cluster purity among all selection strategies, confirming that it promotes structured bias propagation as intended. In contrast, PCoreSet+R shows the highest KLL loss and the lowest cluster purity, which corroborates that maximizing sample diversity in the probability space encourages bias propagation.
A virtuous cycle with few-shot teachers.
Beyond accelerating structured bias propagation, we evaluate the few-shot teacher under ActiveKD. Because acquisition is driven by the student predictions, we can test whether the selected samples by each selection method also benefit the teacher. As shown in the two rightmost panels of Fig.˜7, PCoreSet consistently outperforms competing selection strategies for the teacher. This supports that the structured prediction bias of teacher models propagates to the student: by targeting regions underrepresented in the probability space of student models, PCoreSet also covers underrepresented teacher modes, improving both models and inducing a virtuous cycle of better acquisition and stronger distillation.
@@
Figure 9: The heatmap of output probability vectors from different selection strategies in the first active learning round using the DTD dataset. See Section˜E.1 for results on other datasets.
@@
We analyze the progression of four strategies designed to be maximized throughout AL rounds: 1) uncertainty, 2) class balance, 3) feature space diversity, and 4) probability space diversity (PCoreSet). Criteria 1) and 2) are measured using normalized Shannon entropy, while 3) and 4) are evaluated using normalized average minimum distances. We report the average of each metric across 10 datasets, excluding ImageNet. As shown in Fig.˜8, each method performs best on its respective objective: Entropy maximizes uncertainty, ClassBalanced optimizes the entropy of the class distribution, and Coreset maximizes feature space diversity. As expected, our PCoreSet method achieves the highest diversity in probability space, while also ranking third in uncertainty and class balance, and second in feature diversity. Although Coreset is effective in covering the feature space and offers moderate probabilistic diversity, it performs poorly in class balance, which may be the reason that it underperforms than ClassBalanced baseline.
Figure 10: Samples selected by PCoreSet on Caltech101.
@@
To visualize the sample diversity of PCoreSet, we visualize probability heatmaps for the initial labeled set and for the samples selected after round 1 by each active-learning strategy (Fig.˜9). The heatmaps show that PCoreSet selects samples with highly diverse probability profiles, maximizing distributional coverage relative to the initial set. We also visualize selected images from Caltech101 in Fig.˜10, where PCoreSet chooses visually distinct examples spanning many classes (see Section˜E.2 for more cases). We believe this probabilistic and visual diversity helps us select samples underrepresented in the probability space, thereby more efficiently propagating the teacher’s structured prediction bias during distillation.
@@ 5 Conclusion, Limitation, and Future Work
We introduced ActiveKD, integrating active learning with knowledge distillation by leveraging VLMs’ zero- and few-shot capabilities. We discovered VLMs exhibit structured prediction bias that benefits student learning, and proposed PCoreSet to maximize diversity in probability space when selecting samples. Experiments across 11 datasets show PCoreSet outperforms existing methods, especially when labeled data is scarce, enabling efficient knowledge transfer under limited annotation budgets.
Limitations and future work.
While our experiments focused on VLMs, we believe AL has even greater potential by leveraging growing capabilities of generalist foundation models (Wei et al., 2021; Liu et al., 2023; Bai et al., 2023; Achiam et al., 2023; Team et al., 2023). Given that many task-specific applications face challenges with limited labeled data, leveraging generalist models could offer valuable opportunities for future exploration. Also, our implementation is limited to visual recognition tasks, as they are the most representative in computer vision and provide an ideal testbed. Also, VLMs’ strong zero-shot and few-shot capabilities in visual recognition tasks make them natural candidates for knowledge distillation. However, extending ours to more complex tasks such as object detection and segmentation would be a promising direction with dedicated architectural adaptations.
Reproducibility Statement
We are committed to ensuring the reproducibility of our research. To facilitate replication and extension of our work, we provide comprehensive implementation details throughout the paper and supplementary materials. Specifically, we include: 1) detailed experimental setup in Section˜4.1 and Appendix˜B, covering all hyperparameters, optimization settings, and training procedures; 2) complete algorithmic descriptions of our DHO training framework (Algorithm˜3) and inference procedure (Algorithm˜4); 3) specifications of all model architectures tested, including ResNet18, ResNet50, ViT-T/16, ViT-S/16, and MobileNetV2, along with teacher models CLIP ViT-B/32, CLIP ViT-B/16, and ALIGN; 4) comprehensive dataset information for all 11 benchmarks used in our evaluation (Appendix˜C); and 5) statistical rigor through reporting mean performance with 95% confidence intervals across 5 random seeds for all experiments. Our implementation leverages publicly available pretrained models and standard datasets. We will release our code and experimental configurations upon publication to enable full reproducibility of our results.
Ethics Statement
Our work presents no new ethical concerns as ActiveKD is a purely technical contribution for active learning with knowledge distillation using existing publicly available datasets (ImageNet, Caltech101, StanfordCars, Flowers102, FGVCAircraft, Food101, SUN397, DTD, EuroSAT, UCF101, and OxfordPets) that contain no personally identifiable information. No additional data collection, human subjects research, or sensitive information processing is involved in this work. We acknowledge that vision-language models (CLIP and ALIGN) may contain biases from their pre-training data, which our distillation framework preserves without amplification. The computational requirements are modest (single GPU for training student models across all benchmarks), significantly lower than training vision-language models from scratch, promoting research accessibility while minimizing environmental impact.
@@
  • Theoretical Analysis (Appendix˜A): provides formal analysis of bias propagation in knowledge distillation, including optimal student prediction analysis (Section˜A.1), formal characterization of teacher prediction bias (Section˜A.2), and mathematical proof of how biases transfer from teacher to student models (Proposition˜2).
  • Implementation (Appendix˜B): detailed experimental setup including model configurations, training parameters.
  • Datasets (Appendix˜C): comprehensive description of the datasets used in our experiments, including ImageNet and other benchmarks used for evaluating bias propagation, as well as detailed analysis of class distributions across all datasets.
  • Additional Experiments (Appendix˜D): supplementary experimental results including evaluations on alternative architectures (e.g., MobileNetV2 in Section˜D.1), active learning without knowledge distillation (Section˜D.2), few-shot teacher distillation scenarios (Section˜D.3), and logit adjustment for calibrated teacher prediction (Section˜D.4).
  • Additional Qualitative Results (Appendix˜E): additional results of the proposed method, including heatmap visualizations of selected samples across datasets (Section˜E.1) and qualitative analysis of samples selected by our PCoreSet method (Section˜E.2).
@@ Appendix A Theoretical Analysis
This section presents a formal analysis of bias propagation in knowledge distillation, examining how teacher model biases influence student model predictions. We first analyze the optimal student prediction after distillation from a teacher model in Section A.1, building upon the analytical framework established in DHO (Kang et al., 2025). Subsequently, in Section A.2, we formally characterize the mechanisms through which teacher prediction biases propagate to and constrain the distilled student model, providing the theoretical foundation for our debiasing strategies.
@@ A.1 Optimal Student Prediction after Distillation
Let us first establish our notation. We consider a student model with a feature extractor $h:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that maps inputs to feature representations, and a classification head $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ that maps features to logits, with the final prediction obtained by applying the softmax function $\sigma:\mathbb{R}^{C}\rightarrow\Delta^{C-1}$ to these logits, resulting in $f_{r}(x)=\sigma(g(h(x)))$, where $r$ denotes the active learning round. Following DHO (Kang et al., 2025), we consider two target probability distributions: the ground truth label distribution $y$, typically one-hot encoded vectors where $y_{c}=1$ for the true class $c$ and 0 elsewhere, and the teacher’s softened distribution $f(x)$ for input $x\in\mathcal{X}$.
@@
For the Dual-Head Approach, in the Dual-Head Optimization (DHO) framework (Kang et al., 2025), shared features $h(x)$ are extracted from input $x$ and two specialized classification heads are employed: $g_{\text{CE}}(h(x))$, optimized for ground truth labels using cross-entropy loss, and $g_{\text{KD}}(h(x))$, optimized for teacher predictions using KL divergence. This approach addresses the constrained optimization problem:
@@
The DHO framework (Kang et al., 2025) provides a formal proof that under appropriate optimization conditions, the dual-head inference is equivalent to single-head inference when setting $\alpha=\lambda$ and $\beta=\tau$, meaning that for any choice of distillation weight $\lambda$ and temperature $\tau$, the two approaches converge to the same prediction.
@@ A.2 Formal Definition of Teacher Prediction Bias
Let the teacher model $f$ exhibit structured prediction bias as defined in Definition˜2, with $\{\mu_{k}\}_{k=1}^{K}$ and $\{r_{k}\}_{k=1}^{K}$. Assume the student model $f_{r}$ is trained via KD from $f$ using the loss $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, and satisfies $\sup_{x}\|f_{r}(x)-f^{*}(x)\|_{1}\leq\varepsilon$ for all $x\in\mathcal{X}$, where $f^{*}(x)=\lambda y+(1-\lambda)f(x)$ and $y\in\{0,1\}^{C}$ denotes the one-hot label of $x$. Then student predictions $f_{r}(x)$ also exhibit structured prediction bias as in Definition˜2. Specifically, for every $x\in\mathcal{X}$, there exists $k\in[K]$ such that:
@@
where the propagated centroid is defined as $\hat{\mu}_{k}(x)=\lambda y+(1-\lambda)\mu_{k}$, and the adjusted radius is $\hat{r}_{k}=(1-\lambda)r_{k}+\varepsilon$. Since $y$ is one-hot and $\mu_{k}$ is fixed, the set of possible centers $\{\hat{\mu}_{k}(x)\}$ is finite and contained in $\Delta^{C-1}$, thus satisfying the condition of Definition˜2 with at most $C\cdot K$ clusters.
@@ Appendix B Implementation Details
Algorithm 3 DHO Training with zero-shot CLIP (Radford et al., 2021) teacher
@@
[…]
@@
We employ the DHO (Kang et al., 2025) framework for knowledge distillation in our approach, as it provides a principled and effective mechanism for transferring knowledge from foundation teacher models to student models. DHO’s key advantage lies in its dual-head architecture that maintains separated optimization paths for labeled data and teacher knowledge, thereby preserving the distinct learning signals throughout training. Algorithm˜3 details the DHO training procedure with our zero-shot CLIP teacher, while Algorithm˜4 outlines the inference process that effectively combines both optimization pathways.
@@
Student: DINO self-supervised ResNet-50 (Caron et al., 2021) Teacher: CLIP ResNet-50 (Radford et al., 2021) Active learning rounds: 8 Initial setting: 1-shot (single image per class) Query size: 1,000 Unlabeled pool: 100,000 samples randomly selected with seed Random seeds: 5 different seeds KD parameters: $\zeta=0.01$, $\eta=2$, $\lambda=0.5$ DHO parameters: $\alpha=0.4$, $\beta=0.5$ Validation: No validation split used Epochs: 20 for first 7 rounds, 50 for final round Optimizer: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) Learning rate: $1\times 10^{-3}$ Weight decay: $1\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256)
@@
Student: ImageNet pre-trained ResNet-18 (He et al., 2016), MobileNetV2 (Sandler et al., 2018), and ImageNet-21k (Deng et al., 2009) pre-trained ViT-Tiny (Wu et al., 2022) Teacher: CLIP ResNet-50 (Radford et al., 2021) Active learning rounds: 16 Initial setting: 1-shot (single image per class) Query size: Equal to number of classes per round Unlabeled pool: All training samples except labeled set Random seeds: 5 different seeds KD parameters: $\zeta=0.01$, $\eta=2$, $\lambda=0.5$ DHO parameters: $\alpha=0.5$, $\beta=1$ Validation: No validation split used Epochs: 200 for all rounds Optimizer: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) Learning rate: $1\times 10^{-3}$ Weight decay: $1\times 10^{-2}$ Batch size: 128 (labeled: 64, unlabeled: 64)
@@
Teacher: CLAP (Silva-Rodriguez et al., 2024) on CLIP ResNet-50 (Radford et al., 2021) Setup: Following original CLAP paper (Silva-Rodriguez et al., 2024) Modification: Learning rate reduced from 0.1 to 0.01 for better convergence
@@
Recognizing the potential for improving teacher network performance, particularly in low-data regimes, we explored few-shot learning techniques to enhance the distillation process. After considering various approaches (Jia et al., 2022; Zhou et al., 2022b; a; Khattak et al., 2023a; Zhu et al., 2023; Khattak et al., 2023b; Menghini et al., 2023; Zhao et al., 2024; Roy & Etemad, 2023; Zhang et al., 2024; Lafon et al., 2024), we incorporated CLAP (Silva-Rodriguez et al., 2024), a method that notably does not require a validation set (Silva-Rodriguez et al., 2024; Murugesan et al., 2024; Morales-Álvarez et al., 2024), as our teacher model. For training the teacher, we followed the setting from the original CLAP paper (Silva-Rodriguez et al., 2024), with only one modification: reducing the learning rate from 0.1 to 0.01 for better convergence in our setting. This configuration represents a more realistic scenario wherein both teacher and student networks evolve simultaneously within the active distillation framework, better reflecting practical applications where pre-trained teacher models may not be available.
@@
BADGE (Ash et al., 2019) adopts a hybrid approach to maximize uncertainty and diversity simultaneously, utilizing the gradient of the final linear classifier layer which has $H\times C$ elements, where $H$ is the hidden dimension of the backbone feature and $C$ is the number of classes in the teacher model. However, DHO (Kang et al., 2025) extends beyond conventional single classifier architectures to dual classifiers that follow different optimization paths for labeled ground truth and teacher distillation signals. The dual classifier approach provides complementary information from both learning objectives, enhancing model performance through this specialized learning framework. In order to adopt DHO into the BADGE selection method, we manually calculate the gradients of both classifiers and concatenate them to form a gradient representation with $2\times H\times C$ elements. This concatenated gradient naturally extends the original BADGE selection algorithm while preserving its core k-means++ clustering mechanism.
@@
We implement the Class-Balanced Selection algorithm (Algorithm˜5) as our baseline method to address class distribution bias. The algorithm assigns pseudo-labels to unlabeled samples based on model predictions, analyzes class distribution in the labeled set, and computes inverse weights proportional to each class’s representation—giving higher weights to underrepresented classes. It then allocates the query budget across classes according to these weights and randomly selects samples from each class partition. This approach effectively counteracts class imbalance by prioritizing underrepresented classes.